# A Comparison of Galacticus and COZMIC WDM Subhalo Populations

### v1

@@ 4 Discussion
Other factors that could influence the results for COZMIC subhalo populations include inherent limitations associated with working with cosmological zoom-in simulations. A prominent limiting feature of such simulations is the mass resolution, which imposes a minimum mass at which dark matter subhalos can be resolved. In this work, the halo mass resolution of halos is set at $1.2\times 10^{8}\mathrm{M}_{\odot}$, which corresponds to 300 simulation particles. This resolution was chosen as Nadler et al. (2023) pointed out that halo properties are reasonably well-converged at this limit. Despite this, halo masses around the simulation halo resolution become prone to numerical effects such as artificial tidal disruption (van den Bosch & Ogiya, 2018). Such numerical effects have the strongest influence on lower mass halos, which is the regime where WDM predicts the largest deviation from CDM. It should be noted, however, that there is still ongoing discussion around the impact of artificial tidal disruption and other numerical artifacts in cosmological zoom-in simulations (Benson & Du, 2022; Paun et al., 2025; He et al., 2025).
The Galacticus semi-analytic model also has certain features that could impact the results of its WDM subhalo populations. Subhalo evolution is dominated by non-linear physical processes, such as tidal stripping, tidal heating, and dynamical friction. Galacticus makes analytic approximations to model these phenomena, thereby increasing computational efficiency (Pullen et al., 2014). Previous works have shown that when calibrated against CDM $N$-body simulations, Galacticus is able to accurately reproduce corresponding summary statistics (Yang et al., 2020; Du et al., 2024) despite its analytic approximations.
@@ 5 Conclusions
As a semi-analytic model, Galacticus is more computationally efficient than $N$-body simulations. It has been previously shown that Galacticus is effective in reproducing summary statistics of CDM subhalo populations (Yang et al., 2020; Nadler et al., 2023), and the results from this paper extend the capability of the SAM to warm dark matter models over a wide range of $m_{\text{WDM}}$ masses. While this study provides a valuable first step, comparing Galacticus subhalo populations with future WDM $N$-body simulations will enhance the statistical robustness of our conclusions. Having an accurate population of WDM subhalos generated through a SAM enables more computationally efficient analyses across multiple areas of astrophysics and cosmology, such as results in structure formation, gravitational lensing, observed satellite population statistics, stellar dynamics, and direct detection experiments.

### v2

@@ 4 Discussion
Other factors that could influence the results for COZMIC subhalo populations include inherent limitations associated with working with cosmological zoom-in simulations. A prominent limiting feature of such simulations is the mass resolution, which imposes a minimum mass at which dark matter subhalos can be resolved. In this work, the halo mass resolution of halos is set at $1.2\times 10^{8}\mathrm{M}_{\odot}$, which corresponds to 300 simulation particles. This resolution was chosen as Nadler et al. (2023a) pointed out that halo properties are reasonably well-converged at this limit. Despite this, halo masses around the simulation halo resolution become prone to numerical effects such as artificial tidal disruption (van den Bosch and Ogiya, 2018). Such numerical effects have the strongest influence on lower mass halos, which is the regime where WDM predicts the largest deviation from CDM. It should be noted, however, that there is still ongoing discussion around the impact of artificial tidal disruption and other numerical artifacts in cosmological zoom-in simulations (Benson and Du, 2022; Paun et al., 2025; He et al., 2025).
The Galacticus semi-analytic model also has certain features that could impact the results of its WDM subhalo populations. Subhalo evolution is dominated by non-linear physical processes, such as tidal stripping, tidal heating, and dynamical friction. Galacticus makes analytic approximations to model these phenomena, thereby increasing computational efficiency (Pullen et al., 2014a). Previous works have shown that when calibrated against CDM $N$-body simulations, Galacticus is able to accurately reproduce corresponding summary statistics (Yang et al., 2020a; Du et al., 2024) despite its analytic approximations.
@@ 5 Conclusions
As a semi-analytic model, Galacticus is more computationally efficient than $N$-body simulations. It has been previously shown that Galacticus is effective in reproducing summary statistics of CDM subhalo populations (Yang et al., 2020b; Nadler et al., 2023a), and the results from this paper extend the capability of the SAM to warm dark matter models over a wide range of $m_{\text{WDM}}$ masses. While this study provides a valuable first step, comparing Galacticus subhalo populations with future WDM $N$-body simulations will enhance the statistical robustness of our conclusions. Having an accurate population of WDM subhalos generated through a SAM enables more computationally efficient analyses across multiple areas of astrophysics and cosmology, such as results in structure formation, gravitational lensing, observed satellite population statistics, stellar dynamics, and direct detection experiments.